B4b. What is break away and break in point? How to determine them? (around 70 words) (10 marks)
Answer #1

BREAK POINT The point at which two (08) more poles meet is called break point in the root locus BREAKAWAY POINT » If there exIt happens in case of under damped response, where the poles lies on the Imaginary & real asis [complex poles poles Nuo 70 ZeG(S)H(S) = ķ Example s(s+) Characteristic eq → I+GH=0 Itk sist) s(s+i) tk zo K=-g²-s dk ds -25-1=0 s=-1/2 > Breakaway pomy p